Q. 1 mole of Fe reacts completely with0.65 mole of O2 to give a mixture of only FeO and Fe2O3. Mole ratio of ferrous oxide to ferric oxide is:-

(a)3:2
(b)4:3
(c)20:13
(d)none of these

Grade:11

1 Answers

Chetan Mandayam Nayakar
12 years ago

let ferrous be a moles and b be ferric moles, a+(1.5b)=2*0.65,a+2b=1